OpenStudy (anonymous):

I need help with 4 questions! Plz help!! Ill Medaaall!! 1)What is the slope of the line that passes through the points (3, 4) and (5, –4)?

OpenStudy (anonymous):

#1 is Positive 2

OpenStudy (anonymous):

@Mason_Jones573 But, thats not a choice... A) 4 B) 0 C) -1/4 D) -4

OpenStudy (anonymous):

A

OpenStudy (anonymous):

2) Which is the equation of the line? http://static.k12.com/calms_media/media/243500_244000/243564/1/5870bcf844c8d5714dcd90065954a9c99a2462c5/VHS_PA_S2_02_L208_L307_Q7.gif A. x + 3y = 9 B. 3x – y = 9 C. x – y = 9 D. 3x + y = 9

OpenStudy (anonymous):

Its a medal per question. # 2 is 3x – y = 9

OpenStudy (anonymous):

I can close this one and tag u in the others if u want

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok